OpenStudy (anonymous):

Simplify the expression (0.4)to the third power A)0.64 B)1.2 C)12 D)0.064

OpenStudy (kendricklamar2014):

0.4^3 = 0.064

OpenStudy (unanimoose):

Yep.

OpenStudy (anonymous):

http://video.search.yahoo.com/video/play;_ylt=AwrT6VzLUNlUABsA7nQPxQt.;_ylu=X3oDMTB0ZWVkYm84BHNlYwNzYwRjb2xvA2dxMQR2dGlkA1lIUzAwMl8x?p=Simplify+the+expression+%280.4%29to+the+third+power&tnr=21&vid=F0E010254326906026F3F0E010254326906026F3&l=87&turl=http%3A%2F%2Fts3.mm.bing.net%2Fth%3Fid%3DUN.608006922477502826%26pid%3D15.1&sigi=11r13p9ug&rurl=http%3A%2F%2Fwww.youtube.com%2Fwatch%3Fv%3D_9LUGQWF30Y&sigr=11a7dsv9c&tt=b&tit=How+to+Simplify+a+Number+to+the+3rd+Power+%3A+Math+Measurements&sigt=11t109ppn&back=http%3A%2F%2Fus.yhs4.search.yahoo.com%2Fyhs%2Fsearch%3Fp%3DSimplify%2Bthe%2Bexpression%2B%25280.4%2529to%2Bthe%2Bthird%2Bpower%26type%3Dvst_coinis_15_01_ch%26param1%3D1%26param2%3Dcd%253D2XzuyEtN2Y1L1Qzu0EyE0FyE0CyCtA0Azz0EyDtCtByCyB0DtN0D0Tzu0StCtDzyzytN1L2XzutAtFyBtFtCtFtAtN1L1CzutCyEtBzytDyD1V1TtN1L1G1B1V1N2Y1L1Qzu2StA0AtA0FyEtA0C0DtGzyzytA0BtGtDzy0B0AtG0BtB0E0BtGyC0CyEyBtB0DtDtB0AzyyC0A2QtN1M1F1B2Z1V1N2Y1L1Qzu2S0CtA0CtDyB0F0CyBtGzyyE0DtBtGyEzy0F0EtG0A0D0E0BtG0CyCyDtBzz0B0BtAtA0BtBtC2Q%2526cr%253D650605814%2526ir%253D142905_a%2526elng%253Den%2526elcl%253Dus%2526a%253Dvst_coinis_15_01_ch%2526f%253D4%2526cat%253Dweb%2526ulng%253Den-US%2526sid%253D33397d67488f6ff94e064f632bbe588d%2526stype%253Dvst_coinis_15_01_ch%2526sesid%253D2fb1b011c755377a0dd64ccc22bbc65c%2526csr%253D0%2526ipblock%253D0%2526b%253DIE%2526bv%253D11.0%2526os%253DWindows%252B7%2526cc%253Dus%2526ip%253D68.227.36.80%2526pa%253Dvosteran%26hsimp%3Dyhs-fullyhosted_003%26hspart%3Diry%26ei%3DUTF-8&sigb=1q5etenbs&hspart=iry&hsimp=yhs-fullyhosted_003

OpenStudy (anonymous):

$$ (0.4)^3 = \bigg(\frac{4}{10}\bigg)^3 = \frac{4^3}{10^3} = \frac{64}{1000} = 0.064 $$